Red light with a wavelength of 725 nm shines through a single slit, and forms a pattern on a screen placed 1.25 m away. If the width of the slit is 18.5 micrometers, find the width of the central maximum (a) in degrees and (b) in centimeters.
1
Expert's answer
2019-04-17T11:12:23-0400

wavelength "\lambda = 725\, \text{nm}"

Distance from slit to screen "L = 1.25\, \text{m}"

Width of the slit "d = 18.5 \, \mu\text{m}"

Find the angle for the first minimum:


"d \sin \theta = \lambda"

It is small, therefore:


"\sin \theta = \frac{\lambda}{d} \approx \frac{x}{L} \Rightarrow x = \frac{\lambda L}{d}"

The width of the central maximum :

b)


"\Delta = 2x = \frac{2\lambda L}{d} = 9.8\, \text{cm}"

a)

"2\theta = \frac{2\lambda}{d}\, \text{rad} =\frac{2\lambda}{d} \cdot 180/\pi = 4.5^\circ"